What is the opportunity cost (in civilian output) of a defense buildup that raises military spending from 3.2 to 3.4 percent of a $22 trillion economy?
Answer by math_tutor2020(3817)


1 billion = 10^9
1 trillion = 10^12

22 trillion = 22 * 10^12

3.2% of 22 trillion
0.032*(22 * 10^12)
704 * 10^9
704 billion

3.4% of 22 trillion
0.034*(22 * 10^12)
748 * 10^9
748 billion

The defense spending has gone from $704 billion to $748 billion
This is an increase of 748 - 704 = 44 billion dollars.

This is the opportunity cost as it could have been used in other government programs, pay down the debt, pay off healthcare/pensions, or perhaps returned back to the citizens in the form of tax refunds.

Answer: 44 billion dollars